# ELO-Rated Sequence Rewards: Advancing Reinforcement Learning Models

### v1

@@ Abstract
Reinforcement Learning (RL) is highly dependent on the meticulous design of the reward function. However, accurately assigning rewards to each state-action pair in Long-Term RL (LTRL) challenges is formidable. Consequently, RL agents are predominantly trained with expert guidance. Drawing on the principles of ordinal utility theory from economics, we propose a novel reward estimation algorithm: ELO-Rating based RL (ERRL). This approach is distinguished by two main features. Firstly, it leverages expert preferences over trajectories instead of cardinal rewards (utilities) to compute the ELO rating of each trajectory as its reward. Secondly, a new reward redistribution algorithm is introduced to mitigate training volatility in the absence of a fixed anchor reward. Our method demonstrates superior performance over several leading baselines in long-term scenarios (extending up to 5000 steps), where conventional RL algorithms falter. Furthermore, we conduct a thorough analysis of how expert preferences affect the outcomes.
@@
00footnotetext: This paper has been accepted by DDCLS 2024.
@@ 1 Introduction
Deep Reinforcement Learning (RL) has achieved remarkable progress in various fields, including Atari games [7], Go [2, 5], poker [12], and video gaming [4, 6]. However, real-world challenges frequently involve sparse environmental feedback. For example, a Go game may consist of approximately 200 decision states, yet there is only a signal of reward from the last state. This scarcity of feedback requires the careful design of reward functions for RL algorithms. Furthermore, the vast state spaces in many games hinder the ability of agents to learn effectively through self-exploration alone.
This situation often calls for domain experts to design the reward function [29], or use imitation learning for pre-training to guide the agent to learn effective policies [1, 30]. Nevertheless, reward functions designed by human experts face three main issues: Firstly, there may be some errors or hidden flaws in the reward function, leading RL agents to exploit these errors to obtain unreasonable rewards; Secondly, the design by human experts might cap the potential of RL agents, making the agent difficult to surpass the top human experts; Lastly, addressing large-scale problems typically requires a team of experts. These not only incurs high costs but also tends to produce policies that lack distinctive styles or characteristics due to the need for consensus of experts.
To address these issues, we took inspiration from the ordinal utility in economics, which states that people do not accurately calculate the actual value of each decision when making a decision, but obtain a decision by comparing the pros and cons of each strategy [14]. It is reasonable to assume that a RL agent may not require exact numerical rewards for effective learning. The preferences of different trajectories might help train an equally or more powerful RL agent.
@@
Preference based RL (PBRL) is the embodiment of this idea: the teacher (human expert) provides his/her the preference for the behavior of two agents (or trajectories), and then the agents use this feedback to update their policies respectively (see Figure 1). Compared with designing a reward function for each state-action pair, PBRL can not only greatly reduce the workload of human experts, but it also makes estimated reward less biased and harder to be exploited. Furthermore, it may guide the different styles of the agent through human preferences, so that humans can better understand the reasons why the agent chooses some certain strategies.
However, there are currently three problems with PBRL . First, due to the possibility of experts making mistakes, the update method of PBRL training is extremely unstable; Second, the preference judgment of experts is based on trajectories, but current RL algorithms are more suitable for the dense reward model; Third, PBRL relies solely on an expert’s preference between two trajectories, rather than leveraging the interrelationships among multiple trajectories to enhance learning.
Therefore, this paper proposes a new PBRL baseline algorithm, ELO rating based RL (ERRL). For converting the ordinal reward into cardinal reward, we adopt the ELO ranking system, which is widely used to measure player strength in zero-sum games. Particularly, we consider human experts’ preferences for trajectories as a zero-sum game, in which the preferences of different trajectories can be converted into the ELO rating for model optimization. To address the issue of sparse trajectory feedback, we utilize maximum likelihood estimation, achieving a more stable reward decomposition algorithm compared to least squares estimation. Additionally, we take some engineering optimization methods to increase the algorithm efficiency. In the experiments, the results on the selected Atari games are comparable to the state-of-the-art results in the literature. We further demonstrate the effectiveness of the proposed ERRL algorithm through the ablation experiments.
@@ 2.1 Episodic RL with Trajectory Feedback
In classic RL settings, the research focus in Markov Decision Process (MDP) model, MDP general defined as a tuple:$\langle\mathcal{S},\mathcal{A},T,R\rangle$: where $\mathcal{S}$ and $\mathcal{A}$ denote the spaces of environment states and action set of agent. $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$ and $R(s,a)$ denote the probability of state transition to $s^{\prime}$ and the value feedback when action $a$ is taken on state $s$.
@@
In real-world RL problem, it is difficult for people to get the reward function $R(s,a)$ of the MDP model, we just can get the result of the game. So, in most cases, RL must face the difficulties of trajectory feedback or sparse rewards. A trajectory segment is a sequence of state-action pairs $\sigma\left((s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{k},a_{k})\right)\in(\mathcal{S}\times\mathcal{A})^{k}$, trajectory feedback RL model general defined as a tuple: $\langle\mathcal{S},\mathcal{A},T,\mathcal{R}\rangle$ [17], Unlike the MDP model, the environment only gives an evaluation of the entire trajectory $\mathcal{R}(\sigma)$ at the end of the game.
@@
In order to improve training efficiency, it is necessary to redistribute the entire reward to each state-action pair [15, 17], convert the trajectory feedback RL problem into a standard form of RL problem. Constructing a proxy reward function $\hat{r}$ to estimate reward on state-action pair like MDP model. In practice, we optimize the proxy reward function using the following loss function.
@@ 2.2 Ordinal Utility Theory and Cardinal Utility Theory
Cardinal utility theory is a popular concept in microeconomics during the 19th and early 20th centuries [18]. The basic idea is that utility can be measured and summed up. The base unit of utility is called the utility unit. Utility unit can be expressed numerical e.g. $(1,2,3\dots)$, just as the unit of length can be expressed in meters. The mainstream RL reward function $R(s,a)$ is similar to the definition of Cardinality Utility,
Ordinal utility theory was proposed by Hicks and Allen in 1934 [14], and the main point is that: utility as a psychological phenomenon cannot be measured because it is impossible to find a reasonable unit for measurement utility; They believe that what consumers do in the market is not to measure the utility of goods, but to rank among different goods.
Similarly, in some RL problems, it is very difficult to get an accurate Cardinality Reward, but it is easier to obtain the preferences on different trajectories. For example, there are generally 80$\sim$100 decision points in chess, it is difficult to estimate the numerical value of particular states even professional chess players, but the final match result of chess can be judged by even beginner.
@@ 2.3 Elo Rating System
The ELO rating system  [21] is an evaluation method created by Arpad Elo to measure the level of players in various zero-sum game competitions. It is a recognized authoritative method for evaluating players’ abilities widely in competitive competitions, such as chess, Go, football, basketball, e-sports and other competitive competitions. Each player has a hidden rating score, which represents the player’s average competitive level. The higher the score, the stronger the player’s ability. However, due to the randomness of the game and the status of the players, there often exists a gap between the on-the-spot scores and the hidden scores. At the ELO rating model, the hidden score needs to add a fluctuation error, to represent the real ability of the on-the-spot performance. The purpose of the ELO rating is to estimate the true abilities (hidden score) of the players through the win-lose relationship as accurately as possible . The final ELO rating model is as follows:
Random variable $\xi$ obey Logistic distribution, denote the gap between on-the-spot scores and the hidden scores:
@@
$R_{A}$ and $R_{B}$ denote the current hidden score of Player A and Player B, let $D=R_{A}-R_{B}$, At this time, the probability of player A defeat player B is:
$P(D)$ $=1/2+F(x)\big{|}_{0}^{\frac{D}{2}}$ (5)
$=\frac{1}{1+e^{\frac{D}{2}}}$
In order to facilitate calculation and to show the ability gap between players intuitively, ELO modifies the variance of the Logistic distribution and set the initial score at 1500.
@@ 2.4 Performance Based RL (PBRL)
PBRL [16] is a typical algorithm using ordinal utility in RL, which divides the entire trajectory into many sub-trajectories with fixed length $k$, and using $\sigma^{A}\succ\sigma^{B}$ indicates that human experts prefer sub-trajectory $\sigma^{A}$ to sub-trajectory $\sigma^{B}$.
If human experts judge $\sigma^{A}\succ\sigma^{B}$, then there is a high probability that $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$:
$R(s^{A}_{1},a^{A}_{1})+R(s^{A}_{2},a^{A}_{2})+\dots+R(s^{A}_{k_{A}},a^{A}_{k_{A}})>R(s^{B}_{1},a^{B}_{1})+R(s^{B}_{2},a^{B}_{2})+\dots+R(s^{B}_{k_{B}},a^{B}_{k_{B}})$
So it is reasonable to assume that under the proxy reward function $\hat{r}$, the probability of an experts making $\sigma^{A}\succ\sigma^{B}$ is:
@@
Under this assumption, store expert’s preference results $\left(\sigma^{A},\sigma^{B}\,\mu\right)$ into the database $\mathcal{D}$, where $\mu$ is the result function, if $\mu(A)=1$ means $\sigma^{A}\succ\sigma^{B}$, $\mu(A)=0$ means $\sigma^{B}\succ\sigma^{A}$, according to the preference, use the cross entropy loss to optimize the proxy reward function $\hat{r}$ and train reinforcement learning through the proxy reward function $\hat{r}$:
@@ 3 Methodology
In this section, we introduce our new method, ELO rating based RL. It abandons the numerical reward, sorts the trajectories through ELO rating, and uses a simplified version of the return decomposition method to assign ELO rating to each state-action pair.
3.1 Trail Reward Estimation Via Elo Rating
Given two trajectories $\sigma^{A}$ and $\sigma^{B}$, suppose that in trajectory $\sigma^{A}$ achieves better results than trajectory $\sigma^{B}$ formulating as $\sigma^{A}\succ\sigma^{B}$ (ie. in Chess, player A wins Player B). Denote $\mathcal{R}(\sigma^{X})=\overline{\mathcal{R}(\sigma^{X})}+\epsilon$ as the reward obtained by this trajectory $\sigma^{X}$, where $\overline{\mathcal{R}(\sigma^{X})}$ represents the expected reward of using the trajectory $\sigma^{X}$, $\epsilon$ denote the reward fluctuation. In general, it is very difficult to estimate $\overline{\mathcal{R}(\sigma^{X})}$, either need to consume huge computing resources to fully explore the entire solution space; or hire human experts to design appropriate intermediate reward to guide the agent to explore more efficiently.
Although it is difficult to give the appropriate reward for each trajectory $\mathcal{R}(\sigma^{X})$, it is much simpler to assess the pros and cons of each trajectory. At this time, it is better to use the idea of ordinal utility, using $\mathcal{U}(\sigma)$ to represent the ELO rating of the trajectory $\sigma$, and use $\mathcal{U}(\sigma)$ as the reward of the entire trajectory to guide RL training. When $\sigma^{A}\succ\sigma^{B}$, it can be regarded as player $\sigma^{A}$ defeated player $\sigma^{B}$ ($S_{A}=1$); if the expert cannot judge the pros and cons of the two trajectories, it can be described a draw, $\sigma^{A}\approx\sigma^{B}$ ($\approx$ indicates that the experts believe that the results of the two trajectories are similar), can be regarded as a draw between the two sides ($S_{A}=0.5$) . If $\sigma^{A}\succ\sigma^{B}$ the event $\hat{\mathcal{U}}(\sigma^{A})>\hat{\mathcal{U}}(\sigma^{B})$ will occur with high probability, and the ELO rating $\hat{\mathcal{U}}(\sigma^{A})$ can be updated to:
@@ 3.2 Indirect Redistribution of Long-Term Reward
In value based RL, the value function itself implies a one-step reward estimate $R(s,a)=G(s,a)-\gamma G(s^{\prime},a^{\prime})$, there is no need to redefine a proxy reward function, We use deep neural network $\hat{G}_{\text{elo}}(s,a)$ to estimate long-term ELO rating on the state-action pair $(s,a)$. The score for the entire trajectory $\sigma^{X}$ is denote as $\hat{\mathcal{U}}(\sigma^{X})$:
@@
The long-term reward for each state-action pair is updated as (proved in Appendix A):
@@ 3.3 Combining RL and Elo Rating
At the code implementation level, our algorithm is compatible with mainstream RL algorithms. Our algorithm migrates perfectly to any RL algorithm with only a modification to the reward function. In Algorithm 1, we integrate actor-critic RL with Elo rating.
@@
Predictably, the preferences of experts will have a large impact on the results of ordinal-based RL . In addition, the $K$ and $\eta$ parameters of ELO rating will also have a great impact on the results theoretically. We will discuss this problem in detail in Section 4 .

### v2

@@ Abstract
Reinforcement Learning (RL) heavily relies on the careful design of the reward function. However, accurately assigning rewards to each state-action pair in Long-Term Reinforcement Learning (LTRL) tasks remains a significant challenge. As a result, RL agents are often trained under expert guidance. Inspired by the ordinal utility theory in economics, we propose a novel reward estimation algorithm: ELO-Rating based Reinforcement Learning (ERRL). This approach features two key contributions. First, it uses expert preferences over trajectories rather than cardinal rewards (utilities) to compute the ELO rating of each trajectory as its reward. Second, a new reward redistribution algorithm is introduced to alleviate training instability in the absence of a fixed anchor reward. In long-term scenarios (up to 5000 steps), where traditional RL algorithms struggle, our method outperforms several state-of-the-art baselines. Additionally, we conduct a comprehensive analysis of how expert preferences influence the results.
@@ 1 Introduction
Deep Reinforcement Learning (RL) has witnessed significant advancements across diverse domains, such as Atari games [7], Go [2, 5], poker [12], and video gaming [4, 6]. However, real-world applications often encounter sparse environmental feedback. For instance, in a game of Go, which may involve around $N\approx 200$ decision states, the reward signal is solely available at the final state. This sparsity necessitates meticulous design of reward functions for RL algorithms. Moreover, the expansive state spaces in many games impede agents from effectively learning through self-exploration.
In response, it is common practice to either engage domain experts in reward function design [29] or employ imitation learning for pre-training to guide agents in learning optimal policies [1, 30]. Nevertheless, reward functions crafted by human experts are confronted with three primary challenges: First, potential errors or latent flaws within the reward function can be exploited by RL agents to obtain spurious rewards; Second, expert-designed functions may impose an upper bound on the performance of RL agents, hindering them from surpassing human expertise; Third, tackling large-scale problems usually demands a team of experts, which not only escalates costs but also often results in policies lacking uniqueness due to the requirement for expert consensus.
To address these challenges, we draw inspiration from the ordinal utility theory in economics, which posits that decision-makers do not compute the exact numerical value of each choice, but rather derive decisions by comparing the relative merits of strategies [14]. This motivates the hypothesis that RL agents may not require precise cardinal rewards for effective learning—instead, preferences over trajectories could suffice to train agents with comparable or superior performance.
@@
Preference-based RL (PBRL) operationalizes this concept: a teacher (human expert) provides preferences over the behaviors of two agents (or trajectories), which the agents then use to update their policies (see Figure 1). Compared to designing a reward function for each state-action pair, PBRL not only reduces the workload for human experts but also yields estimated rewards that are less biased and less susceptible to exploitation. Moreover, human preferences can guide agents to exhibit diverse styles, enabling better interpretability of why certain strategies are chosen.
However, current PBRL approaches face three key limitations: First, training instability arises from potential expert errors in preference judgments; Second, while experts provide trajectory-level preferences, most RL algorithms are optimized for dense reward models; Third, PBRL typically relies on pairwise trajectory preferences rather than leveraging inter-trajectory relationships to enhance learning.
To address these, we propose a novel PBRL baseline algorithm, ELO Rating-based RL (ERRL). For converting ordinal preferences to cardinal rewards, we adopt the ELO ranking system, widely used to measure player strength in zero-sum games. Specifically, we model human trajectory preferences as a zero-sum game, transforming preference signals into ELO ratings for model optimization. To handle sparse trajectory feedback, we employ maximum likelihood estimation (MLE), which yields a more stable reward decomposition than least squares methods. We also introduce engineering optimizations to enhance algorithm efficiency. Experiments on selected Atari games demonstrate that ERRL achieves performance comparable to state-of-the-art baselines. Ablation studies further validate the effectiveness of our approach. The code is available on GitHub.
@@ 2.1 Episodic RL with Trajectory Feedback
In classic RL settings, the research focuses on the Markov Decision Process (MDP) model, generally defined as a tuple: $\langle\mathcal{S},\mathcal{A},T,R\rangle$. Here, $\mathcal{S}$ and $\mathcal{A}$ denote the environment state space and agent action set, respectively. $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$ represents the probability of transitioning to state $s^{\prime}$ when action $a$ is taken in state $s$, while $R(s,a)$ denotes the value feedback for this state-action pair.
@@
In real-world RL scenarios, specifying the reward function $R(s,a)$ of the MDP model is often impractical; instead, only terminal outcomes of trajectories are observable. Consequently, RL agents frequently encounter challenges related to sparse or delayed feedback. A trajectory segment is defined as a sequence of state-action pairs $\sigma=\left((s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{k},a_{k})\right)\in(\mathcal{S}\times\mathcal{A})^{k}$. In response, trajectory-feedback RL models are typically formalized as a tuple: $\langle\mathcal{S},\mathcal{A},T,\mathcal{R}\rangle$ [17]. Unlike standard MDPs, the environment provides only a cumulative evaluation $\mathcal{R}(\sigma)$ over the entire trajectory $\sigma$ upon termination.
@@
To enhance training efficiency, it is essential to redistribute the cumulative trajectory reward to individual state-action pairs [15, 17], thereby transforming the trajectory-feedback RL problem into a standard MDP framework. This involves constructing a proxy reward function $\hat{r}(s,a)$ to approximate the state-action value, analogous to the reward function in traditional MDPs. In practice, we optimize this proxy reward function using the following loss function:
@@ 2.2 Ordinal Utility Theory and Cardinal Utility Theory
Cardinal utility theory is a prominent concept in microeconomics during the 19th and early 20th centuries [18]. Its core idea is that utility can be measured and summed, with the basic unit of utility referred to as the utility unit. Utility units can be expressed numerically (e.g., $(1,2,3,\dots)$), analogous to how length is measured in meters. The mainstream RL reward function $R(s,a)$ mirrors the definition of cardinal utility.
Ordinal utility theory, proposed by Hicks and Allen in 1934 [14], posits that utility as a psychological phenomenon cannot be measured, as no reasonable unit for quantifying utility exists. They argue that consumers in the market do not measure the utility of goods but rather rank different goods.
Similarly, in certain RL problems, obtaining an accurate cardinal reward is highly challenging, whereas acquiring preferences over different trajectories is more feasible. For instance, chess typically involves 80–100 decision points, making it difficult even for professional chess players to estimate the numerical value of specific states. However, even beginners can judge the final match result.
@@ 2.3 Elo Rating System
The ELO rating system [21] is an evaluation framework developed by Arpad Elo to measure player performance in zero-sum games. Widely adopted in competitive domains such as chess, Go, football, basketball, and esports, it serves as a standard for assessing player skill levels. Each player is assigned a latent rating score representing their average competitive ability: higher scores indicate stronger performance. Due to game randomness and player variability, observed outcomes may deviate from latent ratings. In the ELO model, this discrepancy is modeled as a fluctuating error term to reconcile actual performance with true ability. The core objective of ELO rating is to estimate player abilities (latent scores) as accurately as possible from win-loss outcomes. The formal ELO rating model is defined as follows:
Let the random variable $\xi$ follow a Logistic distribution, representing the gap between observed and latent scores:
@@
Denote $R_{A}$ and $R_{B}$ as the current latent scores of Players A and B, respectively, and let $D=R_{A}-R_{B}$. The probability that Player A defeats Player B is:
[…]
$=\frac{1}{1+e^{-\frac{D}{2}}}$
To facilitate computation and intuitively reflect ability differences, ELO modifies the Logistic distribution’s variance and sets the initial rating at 1500.
@@ 2.4 Performance Based RL (PBRL)
PBRL [16] is a typical RL algorithm leveraging ordinal utility, which partitions each trajectory into fixed-length $k$ sub-trajectories, with $\sigma^{A}\succ\sigma^{B}$ denoting human experts’ preference for sub-trajectory $\sigma^{A}$ over $\sigma^{B}$.
If experts judge $\sigma^{A}\succ\sigma^{B}$, it implies a high probability that $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$:
$R(s^{A}_{1},a^{A}_{1})+R(s^{A}_{2},a^{A}_{2})+\dots+R(s^{A}_{k_{A}},a^{A}_{k_{A}})$
$>R(s^{B}_{1},a^{B}_{1})+R(s^{B}_{2},a^{B}_{2})+\dots+R(s^{B}_{k_{B}},a^{B}_{k_{B}})$
Thus, under the proxy reward function $\hat{r}$, it is reasonable to assume the probability of an expert preferring $\sigma^{A}\succ\sigma^{B}$ is:
@@
Under this framework, expert preference results $\left(\sigma^{A},\sigma^{B},\mu\right)$ are stored in database $\mathcal{D}$, where $\mu$ denotes the preference indicator: $\mu(A)=1$ indicates $\sigma^{A}\succ\sigma^{B}$, and $\mu(A)=0$ indicates $\sigma^{B}\succ\sigma^{A}$. Leveraging these preferences, we optimize the proxy reward function $\hat{r}$ using cross-entropy loss and train the RL agent via the proxy reward function $\hat{r}$:
@@ 3 Methodology
In this section, we introduce our novel method, ELO-Rating based RL (ERRL), which replaces numerical rewards with trajectory rankings via ELO ratings and employs a simplified return decomposition to assign ELO-based values to state-action pairs.
3.1 Trajectory Reward Estimation via ELO Rating
Given two trajectories $\sigma^{A}$ and $\sigma^{B}$, suppose $\sigma^{A}$ outperforms $\sigma^{B}$ (denoted $\sigma^{A}\succ\sigma^{B}$, e.g., Player A wins in chess). Let the trajectory reward be $\mathcal{R}(\sigma^{X})=\overline{\mathcal{R}(\sigma^{X})}+\epsilon$, where $\overline{\mathcal{R}(\sigma^{X})}$ is the expected reward of trajectory $\sigma^{X}$ and $\epsilon$ models reward fluctuation. Estimating $\overline{\mathcal{R}(\sigma^{X})}$ is challenging: it either requires exhaustive exploration of the solution space (prohibitive in computation) or human experts to design intermediate rewards for efficient guidance.
While assigning precise rewards $\mathcal{R}(\sigma^{X})$ is difficult, comparing trajectories is more feasible. Leveraging ordinal utility theory, we define $\mathcal{U}(\sigma)$ as the ELO rating of trajectory $\sigma$, using $\mathcal{U}(\sigma)$ as the trajectory reward for RL training. When $\sigma^{A}\succ\sigma^{B}$, this corresponds to ”player” $\sigma^{A}$ defeating ”player” $\sigma^{B}$ (with score $S_{A}=1$); if experts cannot distinguish between trajectories ($\sigma^{A}\approx\sigma^{B}$), this is treated as a draw ($S_{A}=0.5$). The event $\sigma^{A}\succ\sigma^{B}$ implies a high probability that $\hat{\mathcal{U}}(\sigma^{A})>\hat{\mathcal{U}}(\sigma^{B})$, allowing the ELO rating $\hat{\mathcal{U}}(\sigma^{A})$ to be updated as:
@@ 3.2 Indirect Redistribution of Long-Term Reward
In value-based RL, the value function inherently encodes a one-step reward estimate $R(s,a)=G(s,a)-\gamma G(s^{\prime},a^{\prime})$, obviating the need for a proxy reward function. We employ a deep neural network $\hat{G}_{\text{elo}}(s,a)$ to estimate the long-term ELO rating for state-action pair $(s,a)$. The score for the entire trajectory $\sigma^{X}$ is denoted as $\hat{\mathcal{U}}(\sigma^{X})$:
@@
The long-term reward for each state-action pair is updated as (proved in Appendix Appendix A):
@@ 3.3 Combining RL and Elo Rating
At the implementation level, our algorithm is compatible with mainstream RL frameworks. It can be seamlessly integrated into any RL algorithm by modifying the reward function alone. In Algorithm 1, we demonstrate this compatibility by integrating Elo rating with an actor-critic RL framework.
@@
2:for $\ell=1,2,\dots$ do
3:     Collect a rollout trajectory $\sigma$ using the current policy, and store trajectory $\sigma$ into the replayer buffer $\mathcal{D}$.
4:     Sample $N$ trajectories $W=\{\sigma^{i}\in\mathcal{D}\}^{N}_{i=1}$ from the replay buffer.
5:     for $i=1,2,\dots,N$ do
6:         Generates a shuffled integer list $Y$ from 1 to $N$.
7:         Use equation 10 to get a new long term Elo rating estimation. $E(\sigma^{i},\sigma^{Y[i]})=1/(1+10^{(\hat{\mathcal{U}}(\sigma^{Y[i]})-\hat{\mathcal{U}}(\sigma^{i}))/\eta})$
8:     end for
9:     Use equation 14 to gradient update on the critic network.
$\mathcal{L}(\hat{G}^{\prime}_{\text{elo}})=\sum_{i=1}^{N}\sum_{(s^{i},a^{i})\in\sigma^{i}}\left(\frac{K(\mu(\sigma^{i})-E(\sigma^{i},\sigma^{Y[i]}))}{k_{i}}\right)$ (15)
10:     Use critic network to update actor network.
11:end for
Notably, expert preferences are expected to significantly influence the outcomes of ordinal-based RL. Theoretically, the ELO rating parameters $K$ and $\eta$ will also impact performance, which we will analyze in detail in Section 4.